How do I graph the line and find the slope of (2,3) and (-6,-3)?
Answer by lwsshak3(11628)

How do I graph the line and find the slope of (2,3) and (-6,-3)?
Standard form of equation for a straight line: y=mx+b, m=slope, b=y-intercept
For given points:
slope=∆y/∆x=(-3-3)/(-6-2)=-6/-8=3/4
equation:
y=(3/4)x+b
solving for b using coordinates of one of the given points (2,3)
3=(3/4)*2+b
3=3/2+b
b=3/2
equation of line:
y=3x/4+3/2
Graphing the line:
Starting from the first point on the line at the y-intercept (3/2), move 3 units up and 4 units to the right (rise/run) to find the second point on the line.